1.2*10^14 /0.5*10^-24=
Whats the rule for dividing a fraction with negative exponets?

Answer by stanbon(75887)

1.2*10^14 /0.5*10^-24=
Whats the rule for dividing a fraction with negative exponets?
----
Comment: The negative in an exponent means "invert".
1/2^-1 = 1/(1/2) = 2
----
3^-1 = 1/3
---------------------
===================
Your Problem:
1.2*10^14 /0.5*10^-24
---
= (1.2/0.5)(10^14/10^-24)
----
= (2.4)(10^14*10^24)
---
= 2.4*10^42
================
Cheers,
Stan H.